# Implicit Neural Representations of Intramyocardial Motion and Strain

### v3

@@ Abstract
Automatic quantification of intramyocardial motion and strain from tagging MRI remains an important but challenging task. We propose a method using implicit neural representations (INRs), conditioned on learned latent codes, to predict continuous left ventricular (LV) displacement — without requiring inference-time optimisation. Evaluated on 452 UK Biobank test cases, our method achieved the best tracking accuracy (2.14 mm RMSE) and the lowest combined error in global circumferential (2.86%) and radial (6.42%) strain compared to three deep learning baselines. In addition, our method is $\sim 380\times$ faster than the most accurate baseline. These results highlight the suitability of INR-based models for accurate and scalable analysis of myocardial strain in large CMR datasets.
Keywords:
Deep Learning Myocardial Strain CMR Tagging
@@ 2.2 Implicit Neural Representations
such that $X^{\prime}=X+u$ is the new location of the material point in $I_{t}$. With implicit neural representations, we can optimise a multi-layer perceptron (MLP) to fit $f$ using discrete training examples $(X,u)$ (supervised) or by warping $I_{0}$ by $u$ and maximising $NCC(I_{t},I_{0}\circ u)$, where NCCNCC is the normalised cross-correlation (self-supervised).

### v4

@@ Abstract
Automatic quantification of intramyocardial motion and strain from tagging MRI remains an important but challenging task. We propose a method using implicit neural representations (INRs), conditioned on learned latent codes, to predict continuous left ventricular (LV) displacement — without requiring inference-time optimisation. Evaluated on 452 UK Biobank test cases, our method achieved the best tracking accuracy (2.14 mm RMSE) and the lowest combined error in global circumferential (2.86%) and radial (6.42%) strain compared to three deep learning baselines. In addition, our method is $\sim 380\times$ faster than the most accurate baseline. These results highlight the suitability of INR-based models for accurate and scalable analysis of myocardial strain in large CMR datasets. www.github.com/andrewjackbell/Displacement-INR
@@ 2.2 Implicit Neural Representations
such that $X^{\prime}=X+u$ is the new location of the material point in $I_{t}$. With implicit neural representations, we can optimise a multi-layer perceptron (MLP) to fit $f$ using discrete training examples $(X,u)$ (supervised) or by warping $I_{0}$ by $u$ and maximising $NCC(I_{t},I_{0}\circ u)$, where $NCC$ is the normalised cross-correlation (self-supervised).